OpenStudy (anonymous):

The equation of a number and five is eight less than the number. what is the number?

OpenStudy (anonymous):

Your question is not very clear....... are you sure its right?

OpenStudy (anonymous):

if its not x+5=x-8 then its 5x=x-8 orrrrrrrrrrrrrrrrrrrrrrrrrrrrrrrrrrrrrrrr

OpenStudy (anonymous):

I think its 5x = x - 8 4x = 8 x = 2 Am I right?

OpenStudy (anonymous):

@ pratu If 5x= x-8 then 4x=-8

OpenStudy (anonymous):

oh yes..... stupid mistake, answer is -2

OpenStudy (anonymous):

thanks for correcting my mistake.

OpenStudy (anonymous):

whis is it? A.6 b.5 c.11.33 d.10 ?